OpenStudy (anonymous):

if xy =4 then dy/dx = ? and please explain it

OpenStudy (anonymous):

Apply product rule with respect to x:

OpenStudy (anonymous):

how do I solve this then? I am new to calculus

sam (.sam.):

differentiate you'll get xy'+y=0

sam (.sam.):

\[\text{Use the product rule, }\frac{d}{dx}(u v)=v \frac{du}{dx}+u \frac{dv}{dx}\text{, where }u=x\text{ and }v=y\]

sam (.sam.):

Basically for product rule I think of like this, xy =4 (copy left)(differentiate right) + (copy right)(differentiate left) where "left" is "x" and right is "y" so when you differentiate, (x)(dy/dx)+(y)(1)=0

sam (.sam.):

This is implicit differentiation, did you learn the basics of differentiation first?

OpenStudy (anonymous):

no, I started today. I was doing limit before it. Can you suggest me good tutorials on differntiation and derivative ..

OpenStudy (anonymous):

its very helpful what you explained here but I need to clear the basic concepts of derivative first to solve this equation.

OpenStudy (anonymous):

thank you very much ..

sam (.sam.):

have you tried khanacademy?